The range of the sample data is what

Listed below are the top 10 annual salaries​ (in millions of​ dollars) of TV personalities. Find the​ range, variance, and standard deviation for the sample data. Given that these are the top 10​ salaries, do we know anything about the variation of salaries of TV personalities in​ general?

39

38

36

29

19

18

15

11

10.7

9.9


1
Expert's answer
2022-02-07T17:57:03-0500

The range for this data is,

"range=maximum \space value-minimum \space value=39-9.9=29.1"


The variance is given as,

"s^2={\sum x^2-{(\sum x)^2\over n}\over n-1}={6345.5-5089.536\over9}=139.5516"


The standard deviation is,

"s=\sqrt{s^2}=\sqrt{139.5516}=11.8132"


No.

The given list is only for the top 10 annual salaries for TV personalities therefore, we do not know anything about the variation of salaries of TV personalities in​ general as this sample is not representative of the whole population.